OpenStudy (anonymous):

What is the domain of the function f(x) equals root 2x-3?

myininaya (myininaya):

square root? cube root? fourth root? ...?

OpenStudy (anonymous):

Under the square root symbol. I can't type it out on my keyboard sorry.

OpenStudy (anonymous):

\[2x-3\geq0\] \[2x\geq3\] \[x\geq\frac{3}{2}\]

myininaya (myininaya):

2x-3>=0 solve for x and you will have your domain

OpenStudy (anonymous):

Thank you both of you!